# thanks: Retired Professor

### v2

@@ III.2.1 Parametric microcanonical inflection point analysis for the 2D Ising Model
Recent studies employing conventional microcanonical inflection point analysis of this model [30, 45, 46, 47] have reported evidence of higher-order phase transitions in addition to the well-known second-order ferromagnetic/paramagnetic transition. Specifically, two additional transitions were identified: a dependent transition occurring above the critical temperature and an independent transition occurring below it, as illustrated in the next paragraph.
Fig. 7 presents the results of a parametric microcanonical inflection point analysis for $L=96$, where the density of states (DOS) was obtained using the exact solution provided by Paul D. Beale [48]. The black circles represent $\gamma(\bar{\beta})$, and the vertical dashed green line indicates the position of its peak, corresponding to the second-order phase transition at $\bar{\beta}_{tr}=0.43863$. The red squares represent $\delta(\bar{\beta})$. The blue dotted-dashed line marks the local minima with positive values, signifying an independent third-order phase transition at $\bar{\beta}_{tr}=0.44673873$. Finally, the double-dotted-dashed magenta line indicates the local minima with negative values, corresponding to a dependent third-order phase transition at $\bar{\beta}_{tr}=0.43515905$. Previous studies have shown that an independent fourth-order transition, defined by eq (21), in small systems evolves into this third-order transition as the system size (L) increases [30]. This behavior was also observed in our analysis.
Figure 7: (Color online) $\gamma(\bar{\beta})$ and $\delta(\bar{\beta})$ for the Ising model with linear system size $L=96$.
@@ III.2.2 Fisher leading zero critical behavior analysis
Fig. 8 depicts the Fisher’s zeros map for the Ising model on the complex $\mathcal{B}$-plane, an alternative representation to the conventional $x$-complex plane [49] or analogous quantities such as $w=2\sinh{(2\beta)}$ [50]. Notably, circles in the $x$-complex plane are mapped onto vertical lines in the $\mathcal{B}$-plane. Furthermore, circles with radii exceeding unity correspond to negative temperatures ($\beta<0$), which are not displayed in our graphs. For this model, negative temperatures mean antiferromagnetic ground states, as the temperature is measured in units of $J|\sigma|^{2}/k_{B}$. Owing to the symmetry of the density of states (DOS) for the Ising model, the magnitudes of the transition temperatures are identical. In Fig. 8, the real part of the leading zero is determined to be $k_{B}{\beta}_{1}\approx 0.43868$. While exhibiting a vertical line pattern, this map deviates from the characteristic pattern observed in first-order transitions due to the non-uniform spacing of the dominant zeros [51, 52]. Considering system sizes up to $L=128$, the Finite-size scaling (FSS) analysis of the leading zeros yields an estimate for the critical temperature with an error of approximately 0.02% compared to the exact value. Moreover, this FSS analysis provides an estimate for the critical exponent $\nu$, as shown in Eq. (9), with a 0.7% error relative to the expected value of $\nu=1$ for the Ising universality class, thereby corroborating the reliability of the analysis.
@@ III.3 XY Model
The Berezinskii-Kosterlitz-Thouless (BKT) transition, exemplified by the XY model, is a topological phase transition driven by the unbinding of vortices at the temperature $T_{BKT}$. It is characterized by the absence of discontinuities or divergences in any finite-order derivative of the free energy. The XY model describes two-dimensional systems, a prototype being a lattice of spins with continuous symmetry. The state of each spin is characterized by an angular variable, $\theta_{i}$, denoting its orientation with respect to a fixed reference axis within the plane. Interactions are confined to nearest neighbors, and the Hamiltonian of the system assumes the following form:
@@
In previous work, we investigated this model on a square lattice of dimension $L\times L$ employing Fisher’s zeros methodology [53, 54, 55]. An example of these mappings on $\mathcal{B}$ plane, instead of the originally complex-$x$ map, is depicted in Fig. 9. Despite the absence of dominant zeros, the BKT transition was identified by examining the internal border of zeros through FSS analysis, where the BKT transition temperature was estimated to be $T_{BKT}=0.704(3)$. Our findings were in complete concordance with theoretical predictions, enabling the classification of the phase transition as belonging to the BKT universality class. It is worth mentioning that the lack of a dominant zero in Fig 9 implies an analytic free energy as a function of temperature near the BKT transition.
@@
In Fig. 10, we present the parametric microcanonical inflection point analysis for this model. Given that finite-order derivatives of the free energy remain finite and continuous, one can anticipate no indication of a transition in these graphs. Panel (a) of the figure displays $\gamma(\bar{\beta})$, where a local maximum with a negative value is observed. However, this feature does not exhibit any discernible scaling behavior, precluding the identification of a critical point, as expected. Panel (b) illustrates $\delta(\bar{\beta})$ which similarly shows no evidence of a higher-order transition. Higher-order derivatives are not displayed due to the substantial error caused by the sensitivity of numerical derivatives to statistical fluctuations.
Figure 10: (Color online) The parametric microcanonical inflection point analysis for the XY-Model. In panel (a) we show $\gamma(\overline{\beta})$ and in panel (b) $\delta(\overline{\beta})$, for linear system size $L=50,\ 60,\ 70,\ 80,\ 90,100,\ \text{and }200$. The vertical red solid line indicates the inverse temperature at the BKT transition for the XY model estimated from the zeros map.
@@ III.4 Zeeman Model
We chose to study the Zeeman Model as an instance of a system that does not exhibit a phase transition at any finite temperature. This model consists of $N$ non-interacting $1/2$-spins particles in a magnetic field, $\vec{B}$. The Hamiltonian can be expressed as
@@
[…]
The partition function can be written as [56]:
@@
[…]
Thus, by inspecting equation (30), $Z=0$ if $e^{2\mathcal{B}B}=-1$. Thus, $2\mathcal{B}B=\pm i(2k-1)\pi$, for $k=1,2,\cdots$. This leads to $\beta_{k}=0$ and
@@
i.e., the Fisher zeros are evenly distributed along the imaginary inverse temperature axis. This analysis explains the presence of a vertical line pattern of zeros near the imaginary axis on the $\mathcal{B}$ maps (which correspond to a unit circle in the complex-$x$ map) for the LJ cluster and XY-model, interpreted as a transition at infinite temperature. However, a perfect alignment with $\beta=0$ is not observed in our results due to the exclusion of positive energy values ($E>0$) from our simulations.
Furthermore, by considering Stirling’s approximation, which states that $\ln{y!}=y\ln{y}-y+\mathcal{O}(\ln{y})$, the parametric curve for $\bar{\gamma}$ can be deduced to be
$\bar{\gamma}(\bar{\beta})\approx-\left[\frac{1}{B}\cosh{\left(\bar{\beta}B\right)}\right]^{2}.$ (33)
@@ IV Conclusions
In this study, we introduce a novel technique for analyzing phase transitions, utilizing microcanonical quantities with inverse temperature, $\bar{\beta}$, as a parameter. By examining the behavior of thermodynamic quantities, such as entropy and its derivatives, as functions of $\bar{\beta}$, we observe that each type of transition has its own characteristic behavior. To introduce the method we have studied several models in well known universality class.
@@
As a perspective, we highlight the potential of the proposed methodology to serve as a refined and effective analytical tool for the classification of weak first-order transitions. The characterization of these transitions presents considerable analytical challenges, as they can be readily misidentified as second-order transitions due to their subtle signatures. Our primary results suggest that the proposed protocol can elucidate the pre-transitional pseudocritical behavior at $\beta^{*}<\beta_{tr}$, observed in weak first-order transitions at $\beta_{tr}$ [57], thereby offering a potential resolution to the aforementioned challenge.
@@ Acknowledgments
This work received public financial support from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brazil, under grant 409719/2023-4.
Author declarations

### v4

@@ III.2.1 Parametric microcanonical inflection point analysis for the 2D Ising model
Recent studies employing conventional microcanonical inflection point analysis of this model [32, 52, 53, 54] have reported evidence of higher-order phase transitions in addition to the well-known second-order ferromagnetic/paramagnetic transition. Specifically, two additional transitions were identified: a dependent transition occurring above the critical temperature and an independent transition occurring below it, as illustrated in the following paragraphs.
Figure 7 displays the results of a parametric microcanonical inflection point analysis applied to the two-dimensional Ising model, for linear system sizes ranging from $L=28$ to $128$. The DOS used in this analysis was obtained from Beale’s exact solution [55].
The top panel of Fig. 7 illustrates the behavior of $N\gamma(\bar{\beta})$. A FSS analysis for the peak position of $\gamma$, denoted as $\gamma_{m}$, is presented in Fig. 8. The left panel of the figure shows both $\ln(|N\gamma_{m}|)$ versus $\ln(L)$ and $\ln(|t|)$ versus $\ln(L)$, where $t=1-T_{m}/T_{c}$ is the reduced temperature at the peak position, $T_{m}$. Comparing the former plot with Eq. (27), the slope of the linear regression yields a critical exponent $\alpha/\nu=0.083(8)$. This value aligns with the 2D Ising universality class, for which $\alpha=0$. Given that $t\propto L^{-\nu}$, the linear regression of the latter plot provides $\nu=1.047(4)$, which is close to the predicted value of $\nu=1$ for the 2D Ising universality class.
The right panel of Fig. 8 plots $T_{m}$ against $L^{-\nu}$, utilizing the previously estimated $\nu$ value. The critical inverse temperature extrapolated from this linear regression is $T_{c}=2.2692(1)$. This value exhibits a deviation of approximately $6.5\times 10^{-4}\%$ from the exact critical value, $T_{c}\approx 2.269185$, which is marked by the horizontal blue dashed line. It is known that deviations from the simple scaling form for small system sizes [56] necessitate logarithm corrections to properly describe the criticality in this limit [57, 58, 59, 60]. However, the estimates for critical exponents and the high accuracy in determining the transition temperature presented here corroborate the reliability of the analysis.
Figure 7: (Color online) $\gamma(\bar{\beta})$ and $\delta(\bar{\beta})$ for the 2D Ising model on the square lattice. The top panel presents $N\gamma(\bar{\beta})$ and the bottom presents $N^{2}\delta(\bar{\beta})$) for systems sizes ranging from $L=28$ up to $L=128$. The red solid vertical line indicates the critical inverse temperature $\beta_{c}=\ln{(1+\sqrt{2})}/2$.
The bottom panel of Fig. 7 illustrates the behavior of $N^{2}\delta(\bar{\beta})$. For system sizes $L=96$ and $128$, we observe positive-valued local minima in $\delta$ within the region where $\bar{\beta}>\beta_{c}$. The appearance of these minima suggests an independent third-order phase transition. This finding aligns with previous research indicating that a fourth-order transition, defined by Eq. (21), observed in smaller systems, evolves into this third-order transition as the system size $L$ increases [32]. Due to limitations in the availability of the DOS for larger system sizes in this study, an FSS analysis could not be performed for this particular transition.
Drawing parallels with the Binder-cumulant analysis [61] and the established use of Lee-Yang zero ratios across different spatial volumes [62], the critical point can be estimated by the intersection of $\delta(\bar{\beta})$ at zero. Furthermore, the inset in the bottom panel of Fig. 7(b) reveals a negative peak across all system sizes investigated. This feature corresponds to an additional dependent third-order transition within the paramagnetic phase. An attempt at FSS analysis for this transition did not exhibit any scaling behavior, consistent with previous microcanonical studies [32], and thus these results are not presented.
@@ III.2.2 Fisher leading zero critical behavior analysis
Figure 9 depicts the Fisher’s zeros map for the Ising model on the complex $\mathcal{B}$-plane, an alternative representation to the conventional $x$-complex plane [63] or analogous quantities such as $w=2\sinh{(2\beta)}$ [64]. Notably, circles in the $x$-complex plane are mapped onto vertical lines in the $\mathcal{B}$-plane. Furthermore, circles with radii exceeding unity correspond to negative temperatures ($\beta<0$), which are not displayed in our graphs. For this model, negative temperatures mean antiferromagnetic ground states, as the temperature is measured in units of $J|\sigma|^{2}/k_{B}$. Owing to the symmetry of DOS for the Ising model, the magnitudes of the transition temperatures are identical. In Fig. 9, the real part of the leading zero is determined to be $k_{B}{\beta}_{1}\approx 0.43868$. While exhibiting a vertical line pattern, this map deviates from the characteristic pattern observed in first-order transitions due to the non-uniform spacing of the dominant zeros [65, 66].
From the FSS analysis, guided by Eq. (9), the imaginary component of the zeros is observed to approach the real axis, with an extrapolated value $\tau_{c}=1.6(5)\times 10^{-4}$. Furthermore, this analysis yields a value of $\nu=0.993$ for the critical exponent. This result is in agreement with the expected value of $\nu=1$. Moueddene et al. recently indicated that critical coefficients derived from the zeros of the partition function can effectively predict critical phenomena while maintaining a reasonably low computational cost [67]. In our analysis, the critical inverse temperature estimated from the leading zeros is $\beta_{c}=0.440589(2)$, which deviates by approximately $0.02\%$ from the exact value - thus corroborating the previous analysis.
@@ III.3 XY Model
The Berezinskii-Kosterlitz-Thouless (BKT) transition, exemplified by the XY model, is a topological phase transition driven by the unbinding of vortices at temperature $T_{BKT}$. It is characterized by the absence of discontinuities or divergences in any finite-order derivative of the free energy. The XY model describes two-dimensional systems, a prototype being a lattice of spins with continuous symmetry. The state of each spin is characterized by an angular variable, $\theta_{i}$, denoting its orientation with respect to a fixed reference axis within the plane. Interactions are confined to nearest neighbors, and the Hamiltonian of the system assumes the following form:
@@
In previous work, we investigated this model on a square lattice of dimension $L\times L$ employing Fisher’s zeros methodology [68, 69, 70]. An example of these mappings on the $\mathcal{B}$ plane, instead of the originally complex-$x$ map, is depicted in Fig. 10. Despite the absence of dominant zeros, the BKT transition was identified by examining the internal border of zeros through FSS analysis, where the BKT transition temperature was estimated to be $T_{BKT}=0.704(3)$. Our findings were in complete concordance with theoretical predictions, enabling the classification of the phase transition as belonging to the BKT universality class  [^5] The raw data ($S\times E$), used in this study, are publicly available on zenodo.org [78].
@@
In Fig. 11, we present the parametric microcanonical inflection point analysis for this model, where five independent simulations were used to estimate the mean values and error bars. Given that finite-order derivatives of the free energy remain finite and continuous, one can anticipate no indication of a transition in these graphs. Panel (a) of the figure displays $L^{2}\gamma(\bar{\beta})$, where a local maximum with a negative value is observed. However, this feature does not exhibit any discernible scaling behavior, see Fig. 12, precluding the identification of a critical point, as expected. Panel (b) illustrates $L^{4}\delta(\bar{\beta})$, which similarly shows no evidence of a higher-order transition. Higher-order derivatives are not displayed due to the substantial error caused by the sensitivity of numerical derivatives to statistical fluctuations.
Figure 11: (Color online) The parametric microcanonical inflection point analysis for the XY model. In panel (a), we show $\gamma(\overline{\beta})$ and in panel (b) $\delta(\overline{\beta})$, for a linear system size $L=50,\ 60,\ 70,\ 80,\ 90,100,\ \text{and }200$. The vertical red solid line indicates the inverse temperature at the BKT transition for the XY model estimated from the zeros map.
Figure 12: Maximum of $N\gamma$ (left panel) and inverse temperature at peak position (right panel) as functions of the logarithm of linear system size for the XY-model.
@@ III.4 Zeeman Model
We chose to study the Zeeman model as an instance of a system that does not exhibit a phase transition at any finite temperature. This model consists of $N$ non interacting $1/2$-spins particles in a magnetic field, $\vec{B}$. The Hamiltonian can be expressed as
@@
where $n$ represents the number of spins aligned with the magnetic field, and consequently, $(N-n)$ is the number of spins anti-aligned with $\vec{B}$. By directly evaluating the Hamiltonian, the energy density can be expressed as $\mathsf{e}=-Bm=-B(2n/N-1)$, where $m=M/N$ is the magnetization per spin.
The partition function can be written as [72]
@@
$=\Big{[}2\cosh(\mathcal{B}B)\Big{]}^{N}.$ (35)
Thus, by inspecting Eq. (34), $Z=0$ if $e^{2\mathcal{B}B}=-1$. Thus, $2\mathcal{B}B=\pm i(2k-1)\pi$, for $k=1,2,\cdots$. This leads to $\beta_{k}=0$ and
@@
i.e., the Fisher zeros are evenly distributed along the imaginary inverse temperature axis. This analysis explains the presence of a vertical line pattern of zeros near the imaginary axis on the $\mathcal{B}$ maps (which correspond to a unit circle in the complex-$x$ map) for the LJ cluster and XY model, interpreted as a transition at infinite temperature. However, a perfect alignment with $\beta=0$ is not observed in our results due to the exclusion of positive energy values ($E>0$) from our simulations.
Furthermore, by considering Stirling’s approximation, which states that $\ln{y!}=y\ln{y}-y+\mathcal{O}(\ln{y})$, the parametric curve for $N{\gamma}$ can be deduced to be
$N{\gamma}(\bar{\beta})\approx-\left[\frac{1}{B}\cosh{\left(\bar{\beta}B\right)}\right]^{2}.$ (37)
@@
It is pertinent to note that the direct substitution of Eq. (37) into Eq. (24) yields the specific heat as:
$c=(\bar{\beta}B)^{2}\operatorname{sech}^{2}{(\bar{\beta}B)}.$ (38)
This expression exhibits a maximum, characteristic of the well-known Schottky anomaly [73, 74, 75], at $\bar{\beta}B\approx 1.2$, which satisfies the condition $\bar{\beta}B=\coth{(\bar{\beta}B)}$. This maximum arises at a temperature where the thermal energy optimally excites transitions between discrete energy states, resulting in a smooth and broad peak in the specific heat, fundamentally distinct from the sharp features associated with phase transitions.
@@ IV Conclusions
In this study, we introduce a technique for analyzing phase transitions, utilizing microcanonical quantities with inverse temperature, $\bar{\beta}$, as a parameter. By examining the behavior of thermodynamic quantities, such as entropy and its derivatives, as functions of $\bar{\beta}$, we observe that each type of transition has its own characteristic behavior. To introduce the method, we have studied several models in well-known universality class.
@@
As a perspective, we highlight the potential of the proposed methodology to serve as a refined and effective analytical tool for the classification of weak first-order transitions. The characterization of these transitions presents considerable analytical challenges, as they can be readily misidentified as second-order transitions due to their subtle signatures. Our primary results suggest that the proposed protocol can elucidate the pre transitional pseudocritical behavior at $\beta^{*}<\beta_{tr}$, observed in weak first-order transitions at $\beta_{tr}$ [76], thereby offering a potential resolution to the aforementioned challenge.
@@ Acknowledgments
We would like to acknowledge helpful conversations with Dr. Michael Bachmann and Dr. Bruno B. Rodrigues.
This work received public financial support from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brazil, under Grant No. 409719/2023-4.
@@
Data Availability
The data that support the findings of this article are openly available on Zenodo [77, 78].